### v2

# Multiwavelength Observations of the Apparently Non-repeating FRB 20250316A

Abstract
The physical origin of fast radio bursts (FRBs) remains uncertain. Although multiwavelength observations have been widely conducted, only Galactic FRB 20200428D is associated with an X-ray burst from the magnetar SGR J1935+2154. Here, we present multiwavelength follow-up observations of the nearby bright FRB 20250316A, including the Five-hundred-meter Aperture Spherical radio Telescope (FAST), Einstein Probe (EP) X-ray mission, Chandra X-ray Observatory, Wide Field Survey Telescope (WFST) and Space Variable Object Monitor/Visible Telescope (SVOM/VT). The 13.08-hour FAST follow-up campaign without pulse detection requires an energy distribution flatter than those of well-known repeating FRBs, suggesting that this burst is likely a one-off event. A prompt EP follow-up and multi-epoch observational campaign totaling 100 ks led to the detection of an X-ray source within the angular resolution of its Follow-up X-ray Telescope (FXT, ). A subsequent Chandra observation revealed this source to be offset by from the FRB position, and established a 0.5-10 keV flux upper limit of at the FRB position, corresponding to at the 40 Mpc distance of the host galaxy NGC 4141. These results set one of the most stringent limits on X-ray emission from a non-repeating FRB, disfavoring ultra-luminous X-ray sources (ULXs) as counterparts of apparently one-off FRBs and offering critical insights into afterglow models. Our study suggests that an arcsecond localization of both the FRB and its potential X-ray counterpart is essential for exploring the X-ray counterpart of an FRB.
show]yeli@pmo.ac.cn show]xlp@nao.cas.cn show]lick@ihep.ac.cn show]zhuww@nao.cas.cn
1 Introduction
Fast Radio Bursts (FRBs) are millisecond-duration extragalactic radio transients (D. R. Lorimer et al., 2007). Among the nearly one thousand known FRB sources, fewer than 10% exhibit repeating bursts (L. G. Spitler et al., 2016; CHIME/FRB Collaboration et al., 2021; Chime/Frb Collaboration et al., 2023a). Although their origin remains unclear, the association between FRB 20200428D and a pair of X-ray pulses from the Milky Way magnetar SGR 1935+2154 (S. Mereghetti et al., 2020; CHIME/FRB Collaboration et al., 2020; C. K. Li et al., 2021) reveals that at least low-luminosity FRBs can be produced by magnetars.
While multiwavelength counterparts of typical extragalactic FRBs have been searched for, no conclusive association has been established (see B. Zhang 2024 for a review). The campaigns include well-studied repeating sources, such as FRB 20121102A (P. Scholz et al., 2016), FRB 20180301A (S. Laha et al., 2022), FRB 20180916B (M. Pilia et al., 2020; P. Scholz et al., 2020), FRB 20190520B (J. Sydnor et al., 2025), FRB 20220912A (A. M. Cook et al., 2024), and FRB 20240114A (F. Eppel et al., 2025). The most stringent limits in X-ray luminosity come from the nearby (3.6 Mpc) repeating FRB 20200120E (A. B. Pearlman et al., 2025), which exclude X-ray counterparts similar to ultraluminous X-ray sources (ULXs) or the Crab Nebula. However, most bursts from FRB 20200120E are fainter than typical FRBs (M. Bhardwaj et al., 2021; F. Kirsten et al., 2022; B. Zhang, 2024), and its multiwavelength signals may also differ from those of typical FRBs. For apparently non-repeating FRBs, constraints are usually much looser, because they rely primarily on telescopes with large fields of view (FOVs) and relatively poor sensitivities, such as AGILE (F. Verrecchia et al., 2021), Konus-Wind (A. Ridnaia et al., 2024), AstroSat-CZTI (G. Waratkar et al., 2025), and Fermi (G. Principe et al., 2023). Although two apparently non-repeating FRBs, FRB 20190608B and FRB 20200430A, constrain X-ray luminosities to be , the observations occurred nearly two years after the bursts, and both sources are distant, with redshifts (T. Eftekhari et al., 2023).
FRB 20250316A was detected by the Canadian Hydrogen Intensity Mapping Experiment (CHIME, M. Ng & CHIME/FRB Collaboration 2025; CHIME/FRB Collaboration et al. 2025) at 08:33:50.859038(3) UTC on March 16, 2025 (topocentric at 400 MHz). The burst is very bright, with a peak flux of kJy and a fluence of kJy ms. High-cadence radio follow-up radio observations detected no pulse (O. S. Ould-Boukattine et al., 2025; CHIME/FRB Collaboration et al., 2025). The initial localization has arcminute uncertainties. Subsequent localization using the CHIME core and its three outriggers refined the position to (RA, Dec) = (12h09m44.319s, +58d50m56.708s), with 1 positional uncertainties of 68 mas (semi-major axis) and 57 mas (semi-minor axis) at a position angle of east of north (S. Andrew & Chime/FRB Collaboration, 2025; CHIME/FRB Collaboration et al., 2025). This FRB is positionally associated with the nearby spiral galaxy NGC 4141, at a distance of Mpc based on the Tully-Fisher relation(J. G. Sorce et al., 2014; R. B. Tully et al., 2016). While an infrared counterpart candidate from James Webb Space Telescope observations was reported (P. K. Blanchard et al., 2025a), no clear optical counterpart was detected in MMT and Gemini observations (CHIME/FRB Collaboration et al., 2025). Assuming a host galaxy distance of 40 Mpc and a central frequency of 600 MHz for CHIME, the isotropic energy and luminosity of the FRB are and , respectively.
Owing to its proximity and brightness, FRB 20250316A provides an excellent opportunity for multiwavelength counterpart searches and repeatability studies. To achieve this, we conducted radio, X-ray, and optical follow-up observations shortly after its discovery was reported. The structure of this paper is as follows: Section 2 describes the observations, and Section 3 presents the data analysis and results. We discuss the physical implications in Section 4, and conclude with a summary in Section 5.
2 Observations
This section summarizes our multiwavelength observations of FRB 20250316A in radio, X-ray, and optical bands. The timeline relative to the FRB is shown in Figure 1.
2.1 Radio observations
| Start Time | MJD | Exposure | Upper limitsa |
| (UTC) | (seconds) | mJy ms | |
| 2025-03-17 16:45:00 | 60751.69792 | 3600 | 15.3 |
| 2025-03-18 15:38:00 | 60752.65139 | 3600 | 15.3 |
| 2025-03-18 16:50:00 | 60752.70139 | 7200 | 15.3 |
| 2025-03-20 17:28:00 | 60754.72778 | 3900 | 15.3 |
| 2025-03-21 15:39:00 | 60755.65208 | 3900 | 15.3 |
| 2025-03-22 16:56:00 | 60756.70556 | 7200 | 15.3 |
| 2025-03-23 17:00:00 | 60757.70833 | 7200 | 15.3 |
| 2025-03-24 14:58:00 | 60758.62361 | 7200 | 15.3 |
a A burst width of 1 ms was assumed. SNR 7.
Following the report of FRB 20250316A (M. Ng & CHIME/FRB Collaboration, 2025), we triggered the Five-hundred-meter Aperture Spherical radio Telescope (FAST) for follow-up monitoring. Eight sessions between 2025 March 17 and 24 were conducted as part of FAST Director Discretion Time and the FAST FRB key science project (L. Qian et al., 2025) in tracking mode. Data from all 19 beams of the FAST L-band receiver (– GHz) were channelized into 4096 frequency channels, sampled at 49.153 s, and recorded in PSRFITS search mode with dual linear polarization signals 8-bit sampled and processed using ROACH 2. A 1 K noise-switched calibration signal was taken before each observation. The total exposure time was 13.08 hour (13 hours and 5 minutes) across all sessions (Table 1).
2.2 X-ray observations
2.2.1 Einstein Probe
The Einstein Probe (EP, W. Yuan et al. 2022) is an X-ray mission launched in January 2024. It carries two instruments: the Wide-field X-ray Telescope (WXT), a monitoring instrument with a field of view of 3600 deg2, and the Follow-up X-ray Telescope (FXT), which provides a large effective area and an angular resolution of . WXT covered the position of FRB 20250316A in the survey mode six times within a 24-hour window centered on the burst, with a total exposure of 2 hours (Figure 1 and Table 2). However, the FRB position was not covered at the burst time. The temporally closest observation started on 2025-03-16 at 07:11:07 UTC and ended at 07:30:49 UTC, about 1 hour before the burst. This was followed by the first post-burst observation, starting on 2025-03-16 at 11:57:59 UTC, about 3.5 hours after the burst, with an exposure time of 1324 s.
Following the discovery of FRB 20250316A, we observed the FRB field using the FXT (Y. Chen et al., 2021) onboard EP. Our campaign consisted of ten target-of-opportunity observations: nine were taken between 32 hours and 11 days post-burst (Figure 1 and Table 2). An additional late-time observation was carried out on 2025 May 11 to investigate the variability of a previously reported X-ray source in the field, EP J120944.2+585060 (H. Sun et al., 2025a, b).
2.2.2 Chandra
We conducted a follow-up observation of the FRB 20250316A field with the High Resolution Camera (HRC) on the Chandra X-ray Observatory, aiming to search for a potential X-ray counterpart and obtain an arcsecond localization of EP J120944.2+585060. The observation started on 27 March 2025 at 21:31:27 UTC, with an exposure of 10.89 ks, under a Director’s Discretionary Time (DDT) proposal (PI: Sun, https://doi.org/10.25574/30871 (catalog Chandra ObsId 30871)). The target coordinate was (RA, Dec) (J2000) = (12:09:41.52, +58:50:30.14), the position of EP J120944.2+585060. We selected HRC due to the soft spectrum of EP J120944.2+585060 and its superior effective area below 1 keV compared to other Chandra instruments. The HRC’s subarcsecond resolution provides an order-of-magnitude improvement in positional accuracy over the EP-FXT.
| Telescope | Start Time | ObsID | Exposure Time | Flux()a | |
| (UTC) | (seconds) | EP J120944.2+585060 | FRB 20250316A | ||
| EP-WXT | 2025-03-16 07:11:07 | 11916648071 | 1182 | - | |
| EP-WXT | 2025-03-16 11:57:59 | 11916648077 | 1324 | - | |
| EP-FXT | 2025-03-17 16:57:29 | 08500000303 | 2994 | - | |
| EP-FXT | 2025-03-18 15:22:38 | 08500000304 | 2982 | - | |
| EP-FXT | 2025-03-21 05:50:01 | 08500000311 | 5804 | - | |
| EP-FXT | 2025-03-21 17:02:17 | 08500000317 | 10850 | - | |
| EP-FXT | 2025-03-22 17:03:26 | 08500000319 | 10480 | - | |
| EP-FXT | 2025-03-23 04:15:56 | 08500000320 | 16970 | - | |
| EP-FXT | 2025-03-23 15:28:24 | 08500000321 | 19558 | - | |
| EP-FXT | 2025-03-25 10:41:53 | 08500000325 | 21823 | - | |
| EP-FXT | 2025-03-27 22:14:34 | 08500000326 | 19607 | - | |
| EP-FXT | 2025-05-11 02:36:50 | 06800000592 | 7391 | - | |
| Chandra/HRC | 2025-03-27 21:32:27 | 30871 | 10890 | ||
a
Notes: The unabsorbed X-ray fluxes are given as 90% upper limits for non-detections at the position of the FRB20250316A. For sources detected within the instrument’s positional uncertainty, we report the measured fluxes with their 1 statistical uncertainties.
2.3 Optical observations
2.3.1 WFST
The Wide Field Survey Telescope (WFST) is a 2.5-meter aperture facility located in Lenghu, China (T. Wang et al., 2023). It features a 6.5 deg2 field of view enabled by its optical design and a mosaic CCD camera. On 2025 March 22, we carried out 21 -band exposures of the FRB 20250316A field, 180 s each, yielding a total integration time of 63 minutes. The start times, exposure durations, and filters of these observations are listed in Table 3 and Table LABEL:tb:opt1img.
| Start Time | MJD | Texp(s) | Band | maglim | |
| WFST | |||||
| single images | |||||
| 2025-03-22 17:04:19 | 60756.71133 | 180(21) | 22.94 - 23.34 | ||
| stacked images | 180*21 | 24.18 | |||
| SVOM/VT | |||||
| single images | |||||
| 2025-03-23 18:22:00 | 60757.76458 | 60(101) | VTB | 21.90 - 22.10 | |
| 60(101) | VTR | 21.76 - 22.00 | |||
| 2025-03-25 04:22:05 | 60759.18131 | 60(97) | VTB | - | |
| 60(97) | VTR | - | |||
| stacked images | 60*101 | VTB | 24.17 | ||
| 60*101 | VTR | 24.04 | |||
2.3.2 SVOM/VT
The Space-based multi-band astronomical Variable Objects Monitor (SVOM) is a China–France mission designed to detect gamma-ray bursts (J. Wei et al., 2016). In addition to gamma-ray instruments—GRM (gamma-rays), ECLAIR (hard X-rays), and MXT (soft X-rays)—SVOM carries a 43-cm Visible Telescope (VT) with a dichroic beam splitter dividing light into two channels: VT_B (0.4–0.65 m) and VT_R (0.65–1 m). SVOM/VT conduected two runs of observations targeting FRB 20250316A, simultaneously acquiring VT_B and VT_R data. The first run began 2025-03-23 18:22:00 UTC, the second on 2025-03-25 04:22:05 UTC. Each exposure was 60 s. After removing images affected by background noise or distorted PSFs, 101 and 97 good-quality frames were obtained in the two runs for each band. Details are listed in Tables 3 and LABEL:tb:opt1img.
3 Data Analysis and Results
3.1 Radio
Single pulses were searched for in the FAST data using PRESTO (S. M. Ransom, 2001; S. M. Ransom et al., 2002, 2003), with a DM range of 140-180 pc cm-3 and a step size of 0.1 pc cm-3. No pulses with S/N7 were detected, which corresponds to a fluence threshold of 15.3 mJy ms for a fiducial burst width of 1 ms and a bandwidth of 400 MHz (Table 1). This fluence limit translates to an energy upper limit of erg, assuming a source distance of 40 Mpc.
3.2 X-ray
3.2.1 Einstein Probe
Using the WXT data reduction pipeline and calibration database (CALDB, H. Cheng et al., 2025), we analyzed the observations obtained 1.5 hours before and 3.5 hours after FRB 20250316A. No significant X-ray sources (above a 5 significance threshold) were detected within the positional uncertainty radius ( 3 arcmin) centered at the FRB location (Figure 2, upper left panel). For these two observations, we extracted the source counts from a circular region with a radius of 9 arcmin at the FRB position, enclosing over 90% of the photons within the focal spot region as determined by the WXT ground calibration, and the background counts from an annulus with inner and outer radii of 18 arcmin and 36 arcmin, respectively, as shown in Figure 2. We derived the 90% confidence count rate upper limits of 0.0057 counts/s and 0.0036 counts/s, respectively, in the 0.5–4 keV range. Assuming an absorbed power-law spectrum with a photon index of 2 and a Galactic absorption column density of , the flux limits are and , respectively, corresponding to and assuming a luminosity distance of 40 Mpc.
In the first FXT observation, which began at 2025-03-17T16:57:29 with an exposure time of 2.9 ks, no X-ray sources were detected within the CHIME error region. In the second observation, starting at 2025-03-18T15:23:50 with the same exposure time of 2.9 ks, FXT detected a weak, uncatalogued X-ray source, designated as EP J120944.2+585060. The position of this source is generally consistent with the first refined localization provided by CHIME (C. Leung & Chime/Frb Collaboration, 2025). The source was observed in all eight subsequent FXT observations. For each observation, we extracted the source using a circular region with a 1-arcmin radius. The background was selected from an annulus with inner and outer radii of 1 arcmin and 3 arcmin, respectively, where we excluded contamination from bright sources. The FXT data were processed using the FXT data analysis software (fxtsoftware v1.20; http://epfxt.ihep.ac.cn/analysis). The process involved particle event identification, pulse invariant conversion, grade calculation and selection (grade 0–12), bad and hot pixel flag and selection of good time intervals using housekeeping files.
The spectrum extracted from the stacked images of all the FXT observations (Figure 2, upper right panel) was fitted in XSPEC (v12.14.1) with an absorbed power-law model. The best-fit photon index is (uncertainties at the confidence level), assuming the same Galactic absorption as previously introduced. When modeling the spectrum with the absorption as a free parameter, we found negligible intrinsic absorption. The photon index remains consistent with that obtained from the fixed-absorption fit.
The EP-FXT’s angular resolution ( 23 arcseconds in half power diameter, HPD) (Y. Chen et al., 2025) and localization uncertainty ( 10 arcseconds) are insufficient to conclusively associate the EP source with FRB 20250316A. Therefore, a Chandra/HRC observation was requested to investigate the potential association of the X-ray source with the FRB.
3.2.2 Chandra
We used the Chandra Interactive Analysis of Observations (CIAO, A. Fruscione et al. 2006; version 4.17) software package and the Chandra Calibration Database (CalDB; version 4.12) to reduce and analyze the Chandra observations. By applying the vtpdetect in CIAO for source detection, we identified an X-ray source (CXOU J120943.7+585051) within the FXT error circle (Figure 2, lower left panel). The source is located at RA (J2000) = 12:09:43.72, Dec (J2000) = +58:50:51.4, with a positional uncertainty of 0.74 arcsec (68% confidence level, statistical + systematic). Employing a circular source region with a radius of 1 arcsec and a background annulus with inner and outer radii of 3 arcsec and 5 arcsec, respectively, we obtained a detection significance of 8 using the Li-Ma formula (T. P. Li & Y. Q. Ma, 1983). The source count rate was derived using the srcflux tool, yielding a measurement of counts/s (1 uncertainties, 0.1–10.0 keV). Assuming an absorbed power-law model with a photon index and a hydrogen column density , we estimated the unabsorbed 0.5–10 keV flux from the count rate using PIMMS (K. Mukai, 1993) using the average spectral response files provided within the tool. The resulting flux is (1 uncertainties), generally consistent with previous EP-FXT measurements, thereby confirming this source as EP J120944.2+585060. The position of this source is offset by 7 arcseconds from the precise localization of FRB 20250316A provided by the full CHIME/FRB Outriggers (FRB Collaboration et al., 2025). The combined positional uncertainty for the two sources is approximately 0.7 arcsecond (derived from the sub-arcsecond uncertainties of both the FRB and Chandra/HRC positions), meaning the observed spatial separation corresponds to nearly 10. This relatively large spatial separation suggests that EP J120944.2+585060 is not physically associated with FRB 20250316A.
We note the presence of another object, CXOU J120946.7+585054, detected at 4 confidence level within 1 arcmin of the FRB, at the coordinates of RA (J2000) = 12:09:46.71, Dec (J2000) = +58:50:54.1 (with a positional uncertainty of 0.93 arcsec, 68% confidence level, statistical + systematic, green circle in the lower left panel of Figure 2). It is away from CXOU J120943.7+585051 and away from the nominal EP J120944.2+585060. It is difficult for FXT to resolve these two objects, as shown in the stacked FXT image (upper right panel of Figure 2). The measured count rate for CXOU J120946.7+585054, obtained with srcflux, is counts/s (1 uncertainties, 0.1–10.0 keV). Assuming an absorbed power-law model with a photon index and the galactic absorption, we estimated the unabsorbed 0.5–10 keV flux from the count rate using PIMMS to be . Thus, the flux of CXOU J120943.7+585051 measured from the FXT observations are contaminated by 30% due to CXOU J120946.7+585054. This contamination may explain why the FXT fluxes are slightly higher than the Chandra flux, although the difference is not statistically significant.
At the position of FRB 20250316A, we derive a 90% upper limit on the unabsorbed 0.5 – 10 keV flux of , assuming an absorbed power-law spectrum with a photon index of 2 and . For a distance of 40 Mpc from the host galaxy, this corresponds to a 0.5-10 keV X-ray luminosity upper limit of .
3.2.3 Properties of CXOU J120943.7+585051
Although the precise location of CXOU J120943.7+585051/EP J120944.2+585060 provided by Chandra rules out its association with FRB 20250316A, the nature of the X-ray source remains of interest. The source was detected by FXT in observations spanning over 50 days. The keV light curve extracted at the Chandra position shows moderate variability (see Table 2 and Figure 3). However, a test against a constant flux model yielded for 8 degrees of freedom (dof), corresponding to a p-value of 0.419. This high p-value indicates that the observed data are consistent with the null hypothesis of no variability, showing no statistically significant deviation from a constant flux. Together, these results indicate that the flux remained stable within uncertainties across the earlier observations. The late-time observation in May further confirms this steady behavior.
With an estimated flux of inferred from the Chandra observation, this source, if associated with the host galaxy at 40 Mpc, corresponds to a luminosity of . This places it within the typical luminosity range for ULXs, which are thought to be powered by accretion onto a compact object (either a stellar-mass black hole, neutron star, or potentially an intermediate-mass black hole) at a luminosity comparable to or above the Eddington luminosity of stellar black holes (see P. Kaaret et al. 2017 for a review). Alternatively, if CXOU J120943.7+585051 is not associated with the FRB’s host galaxy, it could be a foreground Galactic source (such as a cataclysmic variable) or a background extragalactic object like an active galactic nucleus (AGN). Distinguishing between these possibilities would require multiwavelength observations.
3.3 Optical
3.3.1 WFST
WFST images were reduced with standard optical procedures. After extracting objects using SExtractor (E. Bertin & S. Arnouts, 1996), we cross-matched them with Gaia and refined the WCS using SCAMP (E. Bertin, 2006). Photometric zero points were calibrated with Pan-STARRS -band magnitudes (K. C. Chambers et al., 2016). To examine host details, we stacked images using SWarp (E. Bertin et al., 2002) with median values. The stacked image is shown in the lower right panel of Figure 2, with the positions of FRB 20250316A (red), EP J120944.2+585060 (yellow), and CXOU J120943.7+585051 (cyan) overplotted. FRB 20250316A and CXOU J120943.7+585051 lie in two extended regions, likely star-forming regions of the host galaxy. The center of the star-forming region hosting FRB 20250316A was determined at (RA, Dec) = (182.434782∘, 58.848778∘) using the centroid_com function from Photutils. The FRB lies 1.1′′ from this centroid (215 pc at 40 Mpc), consistent with CHIME/FRB Collaboration et al. (2025) and P. K. Blanchard et al. (2025b). To assess systematic uncertainties, we also measured the offset between the FRB and the peak of the star-forming region, yielding 0.8′′ (157 pc).
After this, we searched for possible transient optical counterparts of FRB 20250316A in individual and the stacked image, using the DESI/Legacy Survey (DESI/LS)111https://www.legacysurvey.org/ band image as a template. Individual images were aligned to the DESI/LS stack with reproject (T. Robitaille et al., 2020) in Python, then subtracted using HOTPANTS (A. Becker, 2015) with a kernel size of 1.5 times the WFST FWHM. No significant objects were found at the FRB 20250316A position in individual subtracted images. The upper limits were estimated as , where ZP is the zero point of each image and is the RMS within 90′′ of the FRB in the residual image. The 5 upper limits are listed in Table 3 and Table LABEL:tb:opt1img. An upper limit of mag was estimated from the stacked image using a similar process, limited by the seeing and DESI/LS depth. This is generally consistent with the limits of FRB optical counterpart from MMT telescope ( on 2025 March 23) and Gemini ( on 2025 March 24) (CHIME/FRB Collaboration et al., 2025).
3.3.2 SVOM/VT
We first stacked the images obtained on 2025-03-23 and 2025-03-25 individually, following the standard procedure using IRAF v2.16 (D. Tody, 1986, 1993): (1) Images with abnormally high backgrounds or distorted stellar profiles were excluded. The remaining images underwent bias, dark, and flat-field corrections. (2) Precise astrometric alignment was performed with astrometry.net (D. Lang et al., 2010; P. Joseph & A. Dagore, 2024) relative to a chosen reference frame using common bright stars. (3) The aligned images were then stacked with IRAF.IMCOMBINE task with an average combination method, incorporating sigma-clipping to reject cosmic rays and bad pixels, improving the signal-to-noise ratio of the final stacked images. For both the VTB and VTR bands, we stacked 101 and 97 frames from 2025-03-23 and 2025-03-25, respectively, each with 60 s exposure. We then searched for possible optical counterparts with a similar process to that of WFST. Because SVOM/VT observations were taken on two different days, and used custom-designed filters ( nm for VTB and nm for VTR), we utilized the images from SVOM/VT itself as templates to perform image subtraction. Since signals temporally closer to the burst are expected to be more significant, the last single and the stacked images from 2025-03-25 were used as templates, for single and stacked image subtraction, respectively. The upper limits are summarized in Table 3 and Table LABEL:tb:opt1img.
4 Physical Implications and Discussions
4.1 Probability of Chance Coincidence
The X-ray source CXOU J120943.7+585051 was detected by Chandra within the FXT error circle. Given its relatively large separation of 7 arcseconds from the FRB position, the source is unlikely to be physically associated with FRB 20250316A. To quantify the likelihood of such a chance alignment, we calculated the probability of chance coincidence in this section. It should be noted that alternative origins for CXOU J120943.7+585051, such as a foreground Galactic X-ray source or a background AGN, cannot be fully ruled out.
Firstly, we estimate the probability to have an X-ray source within a circle with a 10′′ radius on average. Based on the X-ray source density (A. Masini et al., 2020) at EP/FXT limiting flux, in keV, the averaged number of X-ray sources within the EP/FXT angular resolution is .
Secondly, the density of X-ray sources within galaxies may be enhanced. Thus, we also calculated the probability of a random, unrelated ULX appearing within 10 arcseconds of the FRB position in a galaxy. A 10-arcsecond angular separation at 40 Mpc corresponds to a projected physical radius of 1.94 kpc. The likelihood of finding an unrelated ULX within this region depends critically on the host galaxy properties and the surface density of ULXs. For the general galaxy population, the average number of ULXs ( ) is reported to be (K. Kovlakas et al., 2020). For a typical galactic disk of radius 15 kpc, the expected number of ULXs in a 1.94 kpc radius circle is , giving a 1.6% chance probability using a Poisson statistics. However, this probability is significantly enhanced in star-forming regions or star-forming galaxies, where ULXs are substantially more prevalent (D. A. Swartz et al., 2011; H. Feng & R. Soria, 2011). In such active environments, the density of ULXs could be higher by a factor of 5 to 10, corresponding to a chance probability of approximately –% to detect an unrelated ULX-like source within 10 arcseconds in a star-forming galaxy.
In addition, for even nearby FRBs, low-mass X-ray binaries (LMXBs) and high-mass X-ray binaries (HMXBs) may contribute to contamination. Using a similar method as above, and including the luminosity functions of LMXBs, HMXBs, and ULXs (S. Wang et al., 2016; M. Gilfanov et al., 2022), we estimate the average number of unrelated LMXBs, HMXBs, and ULXs expected within the angular resolution of EP/FXT (10′′) and Chandra (1′′) for different luminosity distances , as shown in Figure 4. It shows that the chance of contamination from LMXBs, HMXBs, and ULXs is about 1% for EP/FXT and 0.01% for Chandra. Furthermore, given that around 20% FRBs reside in host galaxies with AGN signals (T. Eftekhari et al., 2023; K. Sharma et al., 2024), the contamination from AGNs or intermediate-mass black holes (IMBHs) is then non-negligible, especially for higher redshifts. These considerations highlight the importance of arcsecond-level localization for both FRBs and their potential X-ray counterparts, although EP/FXT candidates remain useful in most cases.
4.2 One-off or Repeating
One of the most critical questions in FRB research is whether all FRBs repeat. As a nearby FRB, FRB 20250316A provides a valuable opportunity to examine this issue. Observationally, it was detected by CHIME after 269.77 hours of exposures over six years (CHIME/FRB Collaboration et al., 2025), corresponding to an FRB Poisson rate of 0.004 per hour with a 1 range of per hour ( Jy ms, corresponding to erg Hz-1). FAST conducted 13.08 hours of follow-up observations over approximately one week (Section 3.1). Assuming a Poisson distribution, this indicates a 3 upper limit of ( mJy ms, corresponding to erg Hz-1). The left panel of Figure 5 presents FRB rates estimated with CHIME (red dot) and the upper limit from FAST (red triangle). For comparison, we overplot the energy functions of repeating FRBs. The energy functions of repeating FRBs are empirically fitted as a power law, with a slope of , spanning a range from to (C. J. Law et al., 2017; K. Gourdji et al., 2019). A flatter high-energy component has sometimes been proposed for repeating FRBs, such as FRB 20121102A (D. Li et al., 2021; D. M. Hewitt et al., 2022), FRB 20200120E (S. B. Zhang et al., 2024), FRB 20201124A (Y.-K. Zhang et al., 2022; F. Kirsten et al., 2024), and FRB 20220912A (O. S. Ould-Boukattine et al., 2024). Their energy functions are shown as blue lines, with the lower index, higher index, and the break spectral energy labeled. The flatter component typically extends over orders of magnitude in spectral energy. A comparison between FRB 20250316A and repeating FRBs reveals two possible scenarios. First, the typical power law with a slope of (red solid line) predicts an FRB rate of above the FAST limiting fluence of , more than five orders of magnitude larger than the FAST 3 upper limit. Second, scaling the energy functions of repeating FRBs to the rate of FRB 20250316A at its spectral energy (blue lines) shows that repeating FRBs predict rates 3–7 orders of magnitude higher than the FAST upper limit.
We note that repeating FRBs are usually clustered, with waiting times often described by Weibull distributions (N. Oppermann et al., 2018; L. C. Oostrum et al., 2020; Chime/Frb Collaboration et al., 2020; K. M. Rajwade et al., 2020; F. Kirsten et al., 2024). The Weibull shape parameter characterizes clustering: bursts are clustered if , with typical values – (N. Oppermann et al., 2018; L. C. Oostrum et al., 2020; R. Luo et al., 2020). Using the simulation method of R. Luo et al. 2020, we obtained FRB rate upper limits of 2.0, 0.7, and 0.5 (at FAST limiting fluence) for , 0.5, and 0.9, respectively—still 2–7 orders of magnitude below predictions from repeating FRB energy functions.
On the other hand, some repeating FRBs have 10 bursts, with FRB rates of bursts per hour at a completeness of Jy ms (Chime/Frb Collaboration et al., 2023b). Their energy functions cannot be examined in detail. In order to compare with them, we estimated the FRB rate of FRB 20250316A by combining observations from multiple telescopes. A cumulative FRB rate with multiple telescopes can be estimated as
where is the energy function of FRBs, iterates over different telescopes, and is the threshold of each telescope. To compare with the CHIME repeating FRBs, we adopt following Chime/Frb Collaboration et al. (2023b). Taking into account the observations of FAST, CHIME, and HyperFlash—a 25–32 m European radio telescope network that observed FRB 20250316A for 120 hours at L-band with a limiting fluence of 15 Jy ms (O. S. Ould-Boukattine et al., 2025; CHIME/FRB Collaboration et al., 2025)—and assuming a slope of , we obtain a combined FRB 20250316A rate of hour-1 (red region in the right panel of Figure 5). This region lies below all CHIME repeating FRBs (blue histogram). Using the flattest slope of well-studied repeating FRBs, (FRB 20200120E, S. B. Zhang et al. 2024), the rate is estimated to be hour-1 (gray region), still below all CHIME repeating FRBs. Thus, FRB 20250316A is different from the currently known repeating FRB sample.
CHIME/FRB Collaboration et al. 2025 explored the repeatility of this burst by a simulation, and concludes that the probability of detecting only one burst like FRB 20250316A in repeating FRBs is more significant than 5. With different methods, we have the same conclusion that FRB 20250316A is highly likely to be a one-off event. We caution, however, that this process assumes the energy function of repeating FRBs is stable, whereas temporal variations are possible (D. Li et al., 2021; S. B. Zhang et al., 2024). In addition, we use only the “gold” sample from Chime/Frb Collaboration et al. (2023b). The “silver” sample, which may be contaminated by chance coincidence, may yield CHIME repeating FRBs with a lower rate. As a result, it is hardly to totally exclude the possibility that FRB 20250316A would be detected in some distant future.
4.3 Constraints on multiwavelength counterparts
Different kinds of X-ray counterparts are expected for FRBs. Prompt X-ray pulses with durations ranging from milliseconds to seconds are anticipated in the magnetar framework. Such counterparts were detected in association with bright radio pulses from the Galactic magnetar SGR 1935+2154 (S. Mereghetti et al., 2020; C. K. Li et al., 2021; C. W. Wang et al., 2022). Our observations do not cover the time of FRB 20250316A itself, so we do not focus on this case in detail.
Persistent X-ray emission is also expected in various models. In the magnetar scenario, outbursts of persistent X-ray emission are usually observed during periods of enhanced bursting and flaring activity (V. M. Kaspi & A. M. Beloborodov, 2017; F. Coti Zelati et al., 2018; N. Rea & D. De Grandis, 2025), as in the case of the X-ray burst from the Galactic magnetar SGR 1935+2154 associated with FRB 20200428D (C. D. Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020). Typical long-term X-ray outbursts of magnetars might last several months to years, reaching a maximum luminosity of erg s-1 (V. M. Kaspi & A. M. Beloborodov, 2017), where most of the luminosity is expected to be released in neutrinos (J. A. Pons & N. Rea, 2012). Given our upper limit of erg s-1, we cannot place meaningful constraints on such persistent magnetar outbursts. However, if FRBs are produced by a very young magnetar with fast rotation, one might expect a large amount of rotational power () that could support an X-ray counterpart in the form of pulsed synchrotron emission or a wind nebula. For example, by assuming: 1) a magnetic field of at least G (required for the magnetar to show large X-bursts related to the FRB radio counterpart, R. Turolla et al. 2015), and 2) an X-ray efficiency of (in line with rotationally powered young pulsars, A. Possenti et al. 2002), our deep upper limit implies a maximum rotational power for the young magnetar of erg s-1 (assuming a NS mass of 1.4 and radius of 12 km). Under these assumptions, the putative young magnetar responsible for FRB 20250316A would have a spin period larger than P and a characteristic age longer than 1 yr, possibly in line with the magnetar population of this star-forming galaxy.
Recent FRB environment studies and possible periodicities imply a possible link to binary systems (K. Ioka & B. Zhang, 2020; F. Y. Wang et al., 2022; R. Anna-Thomas et al., 2023; Y. Li et al., 2025; B. Zhang & R.-C. Hu, 2025). Thus, neutron star binary systems, especially LMXB, HMXBs and ULXs, might also be related to FRBs.
In Figure 6, we compare the persistent X-ray luminosity upper limit of FRB 20250316A with those expected from different neutron star classes, both in isolation and in binary systems. We also overplot upper limits from other deep X-ray observations, including the repeating FRB 20200120E (gray open arrow; A. B. Pearlman et al. 2025), as well as the non-repeating FRBs 20200430A and 20190608B (gray filled arrows; T. Eftekhari et al. 2023). The Chandra X-ray upper limit of FRB 20250316A is fainter than nearly all ULXs, while still compatible with an X-ray binary or a magnetar. It is deeper than previously reported limits for other non-repeating FRBs, but less constraining than the limit for the repeating FRB 20200120E. Although the strong variability observed in some ULXs prevents us from completely ruling out these systems as possible progenitors of FRB 20250316A, this result disfavors a direct connection between non-repeating FRBs and ULXs, in agreement with previous findings for the repeating FRB 20200120E.
On the other hand, if FRBs are accompanied by outflows, multiwavelength afterglows are expected, typically lasting for days (S.-X. Yi et al., 2014). In the right panel of Figure 6, we present the X-ray upper-limit constraints obtained from EP-WXT, EP-FXT, and Chandra, together with the optical upper-limit constraints from SVOM and WFST for FRB 20250316A. Here, the 3 values of EP J120944.2+585060 are adopted as upper limits of the FRB. Based on the X-ray upper-limit constraints, and assuming the source is viewed on-axis (which yields the strongest constraint on the kinetic energy), three possible scenarios can be considered.
Ultra-relativistic outflow: If the fireball decelerates before the first WXT observation (i.e., for ), the WXT point at day provides the most stringent kinetic energy constrain, erg. The first FXT point at day then limits the jet opening angle to , requiring a jet break before day to keep the X-ray flux below the FXT measurements (see blue solid line). The corresponding optical emission (orange solid line) is also consistent with the upper limits from SVOM and WFST, as well as the absence of radio counterparts (T. An et al., 2025).
Moderately wide jet: For a larger jet opening angle of , the jet break occurs near the first FXT observation. In this case, the WXT data no longer provide the strongest constraint. Instead, the first FXT point sets the dominant limit, yielding erg. The corresponding X-ray and optical lightcurves are shown by the blue and orange dashed lines.
Low Lorentz factor outflow: If the fireball decelerates around day (near the FXT points), the inferred Lorentz factor is relatively low, . In this scenario, the FXT upper limits do not constrain the jet kinetic energy. Instead, the optical upper limits from SVOM and WFST dominate, requiring erg. The corresponding X-ray lightcurve is shown as the blue dotted line. If we take the MMT upper limit () and and Gemini ( on 2025 March 24) into account, it requires .
These kinetic energy constraints are far larger than the typical FRB energy budget. Therefore, if a hypothetical FRB afterglow exists, the required radio efficiency, , would have to be . It is obvious that an early deep X-ray observation is important to further constrain the X-ray counterpart predicted by the afterglow model.
5 Summary
In this paper, we present multiwavelength follow-up observations of the nearby bright apparent one-off FRB 20250316A, including radio observations with FAST, X-ray observations with EP and Chandra, and optical observations with WFST and SVOM/VT. The non-detection of pulses with the FAST radio telescope indicates that the predicted rate of well-studied repeating FRBs at the FAST limiting fluence is – orders of magnitude higher than FAST’s 3 upper limit, and that the effective FRB rate at the CHIME limiting fluence is lower than the rates of all known repeating FRBs. These results suggest that this burst is distinct from the current sample of repeating FRBs and is likely a one-off event, supporting the results of CHIME/FRB Collaboration et al. 2025. The EP-WXT obtained the first upper limit of at just 3.5 hours post-burst, followed by the EP-FXT achieving a much deeper limit of at 1.4 days after the FRB. The subsequent Chandra observations established the most constraining 0.5–10 keV flux upper limit of , corresponding to an X-ray luminosity upper limit of , which is lower than those of most ULXs. These results place one of the most stringent limits on persistent X-ray emission from a non-repeating FRB, disfavoring the association of FRBs with ULXs and providing critical constraints for theoretical models of FRB progenitors and emission processes. In addition, the combination of the optical and X-ray upper limits constrain the kinetic energy of afterglow models to .
Based on the chance coincidence probability estimation, we should be cautious about future FRB multiwavelength counterpart searches: arcsecond-level localization of both FRBs and their potential X-ray counterparts is essential for a convincing counterpart search.